# Strichartz estimates for the generalized Zakharov-Kuznetsov equation on ×$\mathbb{R}\times\mathbb{T}$ and applications

### v1

@@ 4.2. Improved LWP for mZK
$I_{f}\lesssim\sum_{\begin{subarray}{c}N\sim N_{1}\sim N_{2}\sim N_{3}\\ N\gg 1\end{subarray}}N^{\frac{1}{2}-2s}\lVert Q_{N}^{(\frac{1}{4})}(I_{x}^{\frac{1}{8}}J^{s}u_{i}I_{x}^{\frac{1}{8}}J^{s}u_{j})I_{x}^{\frac{1}{8}}J^{s}u_{l}\rVert_{L_{tx}^{\frac{4}{3}}L_{y}^{2}}\lVert I_{x}^{\frac{1}{8}}f\rVert_{L_{tx}^{4}L_{y}^{2}}$
in the case $(i,j)\neq(0,2)$. The $L_{tx}^{\frac{4}{3}}L_{y}^{2}$-norm can now be estimated exactly as in case (ii.3.3.2), taking into account that
$\lVert\chi_{\{\lvert q\rvert\lesssim N^{\frac{1}{4}}\}}\rVert_{L_{q}^{1}}\sim N^{\frac{1}{4}},$
while the last factor in the $L_{tx}^{4}L_{y}^{2}$-norm can be directly estimated using (19). In both cases, we ultimately obtain
$I_{f}$ $\lesssim\Bigg{(}\sum_{\begin{subarray}{c}N\sim N_{1}\sim N_{2}\sim N_{3}\\ N\gg 1\end{subarray}}N^{\frac{3}{4}-2s}\Bigg{)}\lVert f\rVert_{X_{0,\frac{3}{8}+}}\prod_{i=1}^{3}\lVert u_{i}\rVert_{X_{s,\frac{3}{8}+}}$
and the final step requires $0<\varepsilon\ll 1$ and
$\frac{3}{4}-2s<0\Leftrightarrow\frac{3}{8}<s,$
as desired. Thus, it remains to analyze what can be achieved in the case
$\begin{split}\lvert 3(\xi_{1}+\xi_{3})^{2}-(q_{1}+q_{3})^{2}\rvert&\ll\lvert\xi_{1}+\xi_{3}\rvert^{\frac{1}{2}}\sim\lvert\xi_{0}\rvert^{\frac{1}{2}},\\ \lvert 3(\xi_{0}+\xi_{1})^{2}-(q_{0}+q_{1})^{2}\rvert&\ll\lvert\xi_{0}+\xi_{1}\rvert^{\frac{1}{2}}\sim\lvert\xi_{0}\rvert^{\frac{1}{2}},\\ \lvert 3(\xi_{0}+\xi_{3})^{2}-(q_{0}+q_{3})^{2}\rvert&\ll\lvert\xi_{0}+\xi_{3}\rvert^{\frac{1}{2}}\sim\lvert\xi_{0}\rvert^{\frac{1}{2}}.\end{split}$
Taking into account the sign distribution of the $\xi_{i}$ and $q_{i}$, we obtain
$\lvert\sqrt{3}(\xi_{1}+\xi_{3})+(q_{1}+q_{3})\rvert\ll\lvert\xi_{0}\rvert^{-\frac{1}{2}}\sim\lvert\xi_{\text{max}}\rvert^{-\frac{1}{2}},$
$\lvert\sqrt{3}(\xi_{0}+\xi_{1})+(q_{0}+q_{1})\rvert\ll\lvert\xi_{0}\rvert^{-\frac{1}{2}}\sim\lvert\xi_{\text{max}}\rvert^{-\frac{1}{2}},$
and
$\lvert\sqrt{3}(\xi_{0}+\xi_{3})+(q_{0}+q_{3})\rvert\ll\lvert\xi_{0}\rvert^{-\frac{1}{2}}\sim\lvert\xi_{\text{max}}\rvert^{-\frac{1}{2}},$
and by adding and subtracting these inequalities, we arrive at
$\begin{split}\lvert\sqrt{3}\xi_{0}+q_{0}\rvert&\ll\lvert\xi_{\text{max}}\rvert^{-\frac{1}{2}},\\ \lvert\sqrt{3}\xi_{1}+q_{1}\rvert&\ll\lvert\xi_{\text{max}}\rvert^{-\frac{1}{2}},\\ \lvert\sqrt{3}\xi_{2}+q_{2}\rvert&\ll\lvert\xi_{\text{max}}\rvert^{-\frac{1}{2}},\\ \lvert\sqrt{3}\xi_{3}+q_{3}\rvert&\ll\lvert\xi_{\text{max}}\rvert^{-\frac{1}{2}}.\end{split}$ (56)
Assuming both $\lvert\xi_{1}+\xi_{2}\rvert\gtrsim\lvert\xi_{\text{max}}\rvert^{\frac{1}{4}}$ and $\lvert\xi_{2}+\xi_{3}\rvert\gtrsim\lvert\xi_{\text{max}}\rvert$, inserting the estimates from (56) into (51) yields
$\max_{i=0}^{3}\langle\sigma_{i}\rangle\gtrsim\lvert R\rvert\gtrsim\lvert\xi_{\text{max}}\rvert^{\frac{3}{2}}.$
In the case $\max_{i=0}^{3}\langle\sigma_{i}\rangle=\langle\sigma_{0}\rangle$, we then have
$\lvert\xi_{0}\rvert\langle(\xi_{0},q_{0})\rangle^{s}\langle\sigma_{0}\rangle^{-\frac{1}{2}+2\varepsilon}\lesssim\prod_{i=1}^{3}\lvert\xi_{i}\rvert^{\frac{1}{6}}\langle(\xi_{i},q_{i})\rangle^{\frac{s}{3}-\frac{1}{12}+},$
and by undoing Plancherel and applying Hölder’s inequality, it follows that
$I_{f}$ $\lesssim\lVert I_{x}^{\frac{1}{6}}J^{\frac{s}{3}-\frac{1}{12}+}u_{1}I_{x}^{\frac{1}{6}}J^{\frac{s}{3}-\frac{1}{12}+}u_{2}I_{x}^{\frac{1}{6}}J^{\frac{s}{3}-\frac{1}{12}+}u_{3}\rVert_{L_{txy}^{2}}\lVert f\rVert_{X_{0,\frac{1}{2}-2\varepsilon}}$
$\leq\Bigg{(}\prod_{i=1}^{3}\lVert I_{x}^{\frac{1}{6}}J^{\frac{s}{3}-\frac{1}{12}+}u_{i}\rVert_{L_{txy}^{6}}\Bigg{)}\lVert f\rVert_{X_{0,\frac{1}{2}-2\varepsilon}}$
$\lesssim\Bigg{(}\prod_{i=1}^{3}\lVert u_{i}\rVert_{X_{\frac{s}{3}+\frac{1}{4}+,\frac{1}{2}+\varepsilon}}\Bigg{)}\lVert f\rVert_{X_{0,\frac{1}{2}-2\varepsilon}}$
$\lesssim\Bigg{(}\prod_{i=1}^{3}\lVert u_{i}\rVert_{X_{s,\frac{1}{2}+\varepsilon}}\Bigg{)}\lVert f\rVert_{X_{0,\frac{1}{2}-2\varepsilon}},$
where in the penultimate step we applied the Airy $L^{6}$-estimate (17) three times. For the final step, we require precisely
$\frac{s}{3}+\frac{1}{4}+\leq s\Leftrightarrow\frac{3}{8}+\leq s,$
and by choosing $\varepsilon>0$ sufficiently small, we obtain the desired trilinear estimate for every $s>\frac{3}{8}$. In the case $\max_{i=0}^{3}\langle\sigma_{i}\rangle\neq\langle\sigma_{0}\rangle$ (WLOG, $\max_{i=0}^{3}\langle\sigma_{i}\rangle=\langle\sigma_{1}\rangle$), we can argue in an analogous way: we have
$\lvert\xi_{0}\rvert\langle(\xi_{0},q_{0})\rangle^{s}\langle\sigma_{1}\rangle^{-\frac{1}{2}-\varepsilon}$ $\lesssim\lvert\xi_{0}\rvert^{\frac{1}{6}}\langle(\xi_{0},q_{0})\rangle^{-\frac{1}{3}}\lvert\xi_{2}\rvert^{\frac{1}{6}}\langle(\xi_{2},q_{2})\rangle^{\frac{1}{6}-\frac{1}{8}}$
$\ \ \ \cdot\lvert\xi_{3}\rvert^{\frac{1}{6}}\langle(\xi_{3},q_{3})\rangle^{\frac{1}{6}-\frac{1}{8}}\langle(\xi_{1},q_{1})\rangle^{s},$
which implies
$I_{f}$ $\lesssim\lVert I_{x}^{\frac{1}{6}}J^{-\frac{1}{3}}fI_{x}^{\frac{1}{6}}J^{\frac{1}{6}-\frac{1}{8}}\tilde{u}_{2}I_{x}^{\frac{1}{6}}J^{\frac{1}{6}-\frac{1}{8}}\tilde{u}_{3}\rVert_{L_{txy}^{2}}\lVert u_{1}\rVert_{X_{s,\frac{1}{2}+\varepsilon}}$
$\leq\lVert I_{x}^{\frac{1}{6}}J^{-\frac{1}{3}}f\rVert_{L_{txy}^{6-}}\lVert I_{x}^{\frac{1}{6}}J^{\frac{1}{6}-\frac{1}{8}}\tilde{u}_{2}\rVert_{L_{txy}^{6+}}\lVert I_{x}^{\frac{1}{6}}J^{\frac{1}{6}-\frac{1}{8}}\tilde{u}_{3}\rVert_{L_{txy}^{6+}}\lVert u_{1}\rVert_{X_{s,\frac{1}{2}+\varepsilon}}$
$\lesssim\lVert f\rVert_{X_{0,\frac{1}{2}-2\varepsilon}}\lVert u_{2}\rVert_{X_{\frac{3}{8}+,\frac{1}{2}+\varepsilon}}\lVert u_{3}\rVert_{X_{\frac{3}{8}+,\frac{1}{2}+\varepsilon}}\lVert u_{1}\rVert_{X_{s,\frac{1}{2}+\varepsilon}},$
where this time, in the penultimate step, we used variants of (17), obtained by interpolation with the trivial estimates (13) and (39), respectively. Thus, by choosing $\varepsilon>0$ sufficiently small, we also obtain the desired estimate (42) in this case for all $s>\frac{3}{8}$. All that remains is the case where, for at least one of the quantities $\lvert\xi_{1}+\xi_{2}\rvert$, $\lvert\xi_{2}+\xi_{3}\rvert$ say, for $\lvert\xi_{1}+\xi_{2}\rvert$ we have $\lvert\xi_{1}+\xi_{2}\rvert\ll\lvert\xi_{\text{max}}\rvert^{\frac{1}{4}}$. In this case, however, it follows directly from (55) that $\lvert q_{1}+q_{2}\rvert\ll\lvert\xi_{\text{max}}\rvert^{\frac{1}{4}}\sim\lvert\xi_{0}\rvert^{\frac{1}{4}}$, and the argument can then proceed exactly as in the case $\lvert q_{i}+q_{j}\rvert\ll\lvert\xi_{0}\rvert^{\frac{1}{4}}$, $(i,j)\in\{(0,2),(2,3),(1,2)\}$. With this, all possible subcases have been addressed, and the proof of Proposition 4.3 is complete. ∎

### v3

@@ 4.2. Improved LWP for mZK
$c_{0}$ $\lesssim_{b}\sum_{\begin{subarray}{c}q\in\mathbb{Z}\\ {\lvert}2q_{1}-q{\rvert}\lesssim N^{\frac{3}{4}}\end{subarray}}\int_{\mathbb{R}}\chi_{\{{\lvert}\sqrt{3}\xi\mp q{\rvert}\lesssim N^{-\frac{5}{6}}\}}\mathrm{d}\xi$
$\lesssim N^{-\frac{5}{6}}\sum_{\begin{subarray}{c}q\in\mathbb{Z}\\ {\lvert}2q_{1}-q{\rvert}\lesssim N^{\frac{3}{4}}\end{subarray}}1$
$\lesssim N^{\frac{3}{4}-\frac{5}{6}}$
$=N^{-\frac{1}{12}},$
and this ultimately establishes the estimate
${\lVert}S_{N}^{(-\frac{5}{6},\frac{3}{4})}(u,v){\rVert}_{L_{txy}^{2}}\lesssim_{b}N^{-\frac{1}{24}}{\lVert}u{\rVert}_{X_{0,b}}{\lVert}v{\rVert}_{X_{0,b}},\quad b>\frac{1}{2}.$ (46)
Bilinear interpolation of (46) with the trivial bound
${\lVert}S_{N}^{(-\frac{5}{6},\frac{3}{4})}(u,v){\rVert}_{L_{txy}^{2}}\lesssim{\lVert}u{\rVert}_{X_{1+,\frac{1}{4}+}}{\lVert}v{\rVert}_{X_{0,\frac{1}{4}+}}$
then gives
${\lVert}S_{N}^{(-\frac{5}{6},\frac{3}{4})}(u,v){\rVert}_{L_{txy}^{2}}\lesssim N^{-\frac{1}{24}+}{\lVert}u{\rVert}_{X_{0+,\frac{1}{2}-}}{\lVert}v{\rVert}_{X_{0,\frac{1}{2}-}},$ (47)
thereby allowing us to resume the calculations above: An application of (46) to the first factor and of (47) to the second factor leads us to
$I_{f}\lesssim N^{\frac{11}{12}-2s+}{\lVert}f{\rVert}_{X_{0,\frac{1}{2}-2\varepsilon}}\prod_{i=1}^{3}{\lVert}u_{i}{\rVert}_{X_{s,\frac{1}{2}+\varepsilon}},$
and a dyadic summation over all $N\sim N_{1}\sim N_{2}\sim N_{3}\gg 1$ finally gives the desired trilinear estimate (42), provided that
$\frac{11}{12}-2s+<0\Leftrightarrow\frac{11}{24}+<s$
is satisfied. By choosing $\varepsilon>0$ sufficiently small, any $s>\frac{11}{24}$ can be achieved and this concludes the discussion of this subcase. With this, all possible cases have been addressed, and the proof of Proposition 4.3 is complete. ∎